# Symmetry in language statistics shapes the geometry of model representations

### v1

@@ Abstract
Although learned representations underlie neural networks’ success, their fundamental properties remain poorly understood. A striking example is the emergence of simple geometric structures in LLM representations: for example, calendar months organize into a circle, years form a smooth one-dimensional manifold, and cities’ latitudes and longitudes can be decoded by a linear probe. We show that the statistics of language exhibit a translation symmetry—e.g., the co-occurrence probability of two months depends only on the time interval between them—and we prove that the latter governs the aforementioned geometric structures in high-dimensional word embedding models. Moreover, we find that these structures persist even when the co-occurrence statistics are strongly perturbed (for example, by removing all sentences in which two months appear together) and at moderate embedding dimension. We show that this robustness naturally emerges if the co-occurrence statistics are collectively controlled by an underlying continuous latent variable. We empirically validate this theoretical framework in word embedding models, text embedding models, and large language models.
@@ 1 Introduction
In this work, we provide such an organizing principle: that representational geometry reflects pairwise co-occurrence statistics between words. In particular, we give evidence that symmetries in the statistics collectively drive the formation of representational manifolds, allowing us to give a detailed mathematical explanation of the following phenomena:
  • Circles in representation space for cyclical concepts. (engels2024not; modell2025origins) find that LLMs encode concepts such as days of the week, months of the year, or hues on the color wheel with circular geometry. (prieto2025correlations) find a similar structure in simple auto-encoders. (gromov2023grokking) and related work suggest that circular representations may enable LLMs to efficiently compute modular addition tasks (“Five months after November is:”).
@@
  • Empirical evidence that words embodying a continuous latent concept have pairwise co-occurrence statistics that exhibit translation symmetry (Figures 6 and 7).
@@ 3 Co-occurrence model with symmetry
3.1 states that the co-occurrence probability of any two words in $\mathcal{S}$ depends only on their distance on the semantic continuum. This translation symmetry is a strong constraint on the co-occurrence statistics; nonetheless, we show in Figures 6 and 7 that it accurately describes real data. We assume ${\bm{M}}^{\star}\succeq\bm{0}$ simply for ease of presentation; in Section B.3 we relax this assumption and show that our predictions still hold. [^2] In Section 4, we show that the continuum coordinates constitute a continuous latent variable that underlies and modulates the co-occurrence statistics. In this model, the symmetry condition hold by construction, not by assumption.
@@ 3.2 Predicting embedding geometry for $D=1$ lattices
We leverage Figure 2 to establish the parametric curve for the learned feature manifold corresponding to a periodic $D=1$ concept whose statistics are governed by an exponential co-occurrence kernel. This choice is both analytically tractable and often models the true co-occurrence statistics surprisingly well, encompassing cases such as calendar time and historical time (see Figures 6 and 7).
@@ A.1 Main text figures
(Top row.) For our theoretical model, representation vectors are obtained by measuring the corpus statistics, constructing the matrix ${\bm{M}}^{\star}_{\mathcal{S}}$, and using it to estimate the free parameters of the exponential kernel (see Figures 6 and 7). (Middle row.) For static word embeddings, the representation of “May” is perturbed due to its other meaning (i.e., to express possibility/permission/hope). To eliminate the effect of this polysemy, we exclude May from the computation of the PCA basis. (Bottom row.) For LLM internal representations, we use the prompt formulas given in Section A.3. The effect of polysemy on the representation of “May” is absent since the model is able to disambiguate from context (see Figure 13). Since Gemma 2 2B tokenizes digits separately, prompts for historical years that share the last digit or last two digits (e.g., 1735, 1835, and 1935) produce similar internal representations. This explains the bright off-diagonals in the Gram matrix.
@@ B.2 Asymmetric word embeddings from $\bm{M}$$^{\star}$
B.1 implies that both $|{\bm{M}}^{\star}|_{\mathcal{S}}$ and ${\bm{M}}^{\star}_{\mathcal{S}}$ have a translation symmetry (with different convolution kernels). Note that B.1 strictly softens 3.1, which corresponds to the special case ${\bm{M}}^{-}=\bm{0}$. In addition, B.1 is well-supported in both the empirical data (Figures 6 and 7) and in the collective model of Section 4.1. In particular, if the function $g$ is large enough in magnitude, the non-linear effects associated with the logarithm in Equation 15 can lead to both positive and negative values $\mu_{k}$.
@@ Appendix C Additional empirical evidence
In this appendix we compile additional plots that serve as further empirical evidence.
@@
Figure 6: Translation symmetry in month co-occurrence statistics. The four main quadrants of the plot correspond to the matrices ${\bm{M}}^{\star}={\bm{M}}^{+}-{\bm{M}}^{-}$, $|{\bm{M}}^{\star}|={\bm{M}}^{+}+{\bm{M}}^{-}$, $\bm{M}$$^{+}$, and $\bm{M}$$^{-}$, as defined in Equation 25. For each, on the left we display the submatrix of calendar months and observe the circulant structure indicative of time translation symmetry. On the right, we plot the normalized co-occurrence as a function of the time interval between months, averaged over all months. We find excellent agreement with the periodized exponential kernel, except along the main diagonal. We argue that this discrepancy does not affect eigenvector ordering, since adding $\kappa{\bm{I}}$ to any matrix $\bm{A}$} only shifts each eigenvalue by a constant $\kappa$ (leaving the eigengaps unchanged). We use the periodized kernel $C(\Delta x)=\sum_{n\in\mathbb{Z}}\exp(-|\Delta x+2n|/\sigma)$ because, e.g., a co-occurrence of April and January may refer not only to the closest January to April, but also with the following January, and with the previous year’s January, etc.
@@
Figure 11: Embedding geometry for 2-dimensional lattices. We show the embedding geometry for 2D lattices with open BC and exponential kernel. The only difference between the left two and right two plots is that the lattice spacing is smaller on the right ($4\times 4$ grid vs. $31\times 31$ grid). For the denser grid (closer to the continuum limit), the cross sections of the embedding manifold have the same horseshoe-like geometry seen in the 1-dimensional case. In both cases, the center of the grid is faithfully reproduced in representation space, while the edges are distorted. This matches the result reported in (park2025iclr) (e.g., their Fig 1); we hypothesize that LLMs solve their in-context learning task (random walks on a grid) by computing co-occurrence statistics in-context using a matrix-factorization-like mechanism.

### v2

@@ Abstract
The internal representations learned by language models consistently exhibit striking geometric structure: calendar months organize into a circle, historical years form a smooth one-dimensional manifold, and cities’ latitudes and longitudes can be decoded using a linear probe. To explain this neural code, we first show that language statistics exhibit translation symmetry (for example, the frequency with which any two months co-occur in text depends only on the time interval between them). We prove that this symmetry governs these geometric structures in high-dimensional word embedding models, and we analytically derive the manifold geometry of word representations. These predictions empirically match large text embedding models and large language models. Moreover, the representational geometry persists at moderate embedding dimension even when the relevant statistics are perturbed (e.g., by removing all sentences in which two months co-occur). We prove that this robustness emerges naturally when the co-occurrence statistics are controlled by an underlying latent variable. These results suggest that representational manifolds have a universal origin: symmetry in the statistics of natural data.
@@ 1 Introduction
In this work, we provide such an organizing principle: that representational geometry reflects pairwise co-occurrence statistics between words. In particular, we give evidence that symmetry governs these statistics and drives the formation of representational manifolds. This hypothesis underpins our theory and explains the following phenomena:
  • Circles in representation space for cyclical concepts. (engels2024not; modell2025origins) find that LLMs encode concepts such as days of the week, months of the year, or hues on the color wheel with circular geometry. (gromov2023grokking) and related work suggest that circular representations may enable LLMs to efficiently compute modular addition tasks (“Five months after November is:”).
@@
  • Empirical evidence that words embodying a continuous latent concept have pairwise co-occurrence statistics that exhibit translation symmetry (Figures 5 and 6).
@@ 3 Co-occurrence model with symmetry
3.1 states that the co-occurrence probability of any two words in $\mathcal{S}$ depends only on their distance on the semantic continuum. This translation symmetry is a strong constraint on the co-occurrence statistics; nonetheless, we show in Figures 5 and 6 that it accurately describes real data. We assume ${\bm{M}}^{\star}\succeq\bm{0}$ simply for ease of presentation; in Section B.3 we relax this assumption and show that our predictions still hold. [^2] In Section 4, we show that the continuum coordinates constitute a continuous latent variable that underlies and modulates the co-occurrence statistics. In this model, the symmetry condition hold by construction, not by assumption.
@@ 3.2 Predicting embedding geometry for $D=1$ lattices
We leverage Figure 2 to establish the parametric curve for the learned feature manifold corresponding to a periodic $D=1$ concept whose statistics are governed by an exponential co-occurrence kernel. This choice is both analytically tractable and often models the true co-occurrence statistics surprisingly well, encompassing cases such as calendar time and historical time (see Figures 5 and 6).
@@ A.1 Main text figures
(Top row.) For our theoretical model, representation vectors are obtained by measuring the corpus statistics, constructing the matrix ${\bm{M}}^{\star}_{\mathcal{S}}$, and using it to estimate the free parameters of the exponential kernel (see Figures 5 and 6). (Middle row.) For static word embeddings, the representation of “May” is perturbed due to its other meaning (i.e., to express possibility/permission/hope). To eliminate the effect of this polysemy, we exclude May from the computation of the PCA basis. (Bottom row.) For LLM internal representations, we use the prompt formulas given in Section A.3. The effect of polysemy on the representation of “May” is absent since the model is able to disambiguate from context (see Figure 13). Since Gemma 2 2B tokenizes digits separately, prompts for historical years that share the last digit or last two digits (e.g., 1735, 1835, and 1935) produce similar internal representations. This explains the bright off-diagonals in the Gram matrix.
@@ B.2 Asymmetric word embeddings from $\bm{M}$$^{\star}$
B.1 implies that both $|{\bm{M}}^{\star}|_{\mathcal{S}}$ and ${\bm{M}}^{\star}_{\mathcal{S}}$ have a translation symmetry (with different convolution kernels). Note that B.1 strictly softens 3.1, which corresponds to the special case ${\bm{M}}^{-}=\bm{0}$. In addition, B.1 is well-supported in both the empirical data (Figures 5 and 6) and in the collective model of Section 4.1. In particular, if the function $g$ is large enough in magnitude, the non-linear effects associated with the logarithm in Equation 15 can lead to both positive and negative values $\mu_{k}$.
@@ Appendix C Additional empirical evidence
Figure 5: Translation symmetry in month co-occurrence statistics. The four main quadrants of the plot correspond to the matrices ${\bm{M}}^{\star}={\bm{M}}^{+}-{\bm{M}}^{-}$, $|{\bm{M}}^{\star}|={\bm{M}}^{+}+{\bm{M}}^{-}$, $\bm{M}$$^{+}$, and $\bm{M}$$^{-}$, as defined in Equation 25. For each, on the left we display the submatrix of calendar months and observe the circulant structure indicative of time translation symmetry. On the right, we plot the normalized co-occurrence as a function of the time interval between months, with bars indicating one standard deviation of variation across months. We find excellent agreement with the periodized exponential kernel, except along the main diagonal. We argue that this discrepancy does not affect eigenvector ordering, since adding $\kappa{\bm{I}}$ to any matrix $\bm{A}$} only shifts each eigenvalue by a constant $\kappa$ (leaving the eigengaps unchanged). We use the periodized kernel $C(\Delta x)=\sum_{n\in\mathbb{Z}}\exp(-|\Delta x+2n|/\sigma)$ because, e.g., a co-occurrence of April and January may refer not only to the closest January to April, but also with the following January, and with the previous year’s January, etc.
@@
Figure 9: Embedding geometry for 2-dimensional lattices. We show the embedding geometry for 2D lattices with open BC and exponential kernel. The only difference between the left two and right two plots is that the lattice spacing is smaller on the right ($4\times 4$ grid vs. $31\times 31$ grid). For the denser grid (closer to the continuum limit), the cross sections of the embedding manifold have the same horseshoe-like geometry seen in the 1-dimensional case. In both cases, the center of the grid is faithfully reproduced in representation space, while the edges are distorted. This matches the results reported in (park2025iclr) and (noroozizadeh2025deep); we hypothesize that LLMs solve their in-context learning task (random walks on a grid) by computing co-occurrence statistics in-context using a matrix-factorization-like mechanism. This hypothesis is supported by a large literature explaining how transformers learn Markov models by computing the latent variables in-context (bietti2023birth; edelman2024evolution; nichani2024transformers; chen2024unveiling).